One leg of a right triangle is 2 ft longer than the shorter side. If the
length of the hypotenuse is 18 ft, how long is each leg?

Answer by scott8148(5896)

let x="shorter side", so x+2="longer side"

by Pythagoras, x^2+(x+2)^2=18^2 ___ 2x^2+4x-320=0

x^2+2x-160=0 ___ use the quadratic formula to find x